OpenStudy (anonymous):

40 liters of milk contain 10% water.find the quantity of water to added to the milk so that it may contain 20% water.

OpenStudy (anonymous):

10% water means 4 L of water. Therefore, .2(40 + x) = 4 + x X is the amount of water added. The LHS shows 20% of the total amount with the water added; the RHS shows how much water that would be. Hint: multiply both sides by 10 and that means no more decimal. You might find that easier to solve.

OpenStudy (anonymous):

the answer that i get from .2(40 + x) = 4 + x equation is right . but i still not understand how you wrote it. can you please explain it . please. @Boilermaker

OpenStudy (anonymous):

Think about it this way: before you add the water, we can determine the amount of water in the original milk like this... .1(40) = 4 This is because (Percent of water in milk)(volume of milk) = (volume of water) The LHS shows 10% of 40 and the RHS is the amount of water, 4L. Since we have added x L of water, we have increased the total volume by x L, but also the total volume of water by x L. We also want the new mixture to be 20% water. Therefore, .2(40 + x) = 4 + x